OpenStudy (anonymous):

find the points of intersection of the graph of y=-lxl and he circle x^2+y^2=.

OpenStudy (anonymous):

what is the radius of the circle

OpenStudy (anonymous):

the graph y=-|x| where x will always be positive within the absolute value so you can write it such as y=-(x) = -x

OpenStudy (anonymous):

however i don't have the full circle equation so i can't tell you where they intersect

OpenStudy (anonymous):

\[x^2+y^2=r^2\] \[y=\sqrt{r^2-x^2}\] \[\sqrt{r^2-x^2}=-x\]

OpenStudy (anonymous):

\[r^2-x^2=(-x)^2=x^2\] \[r^2=2x^2\] \[\frac{r^2}{2}=x^2\] \[\frac{\sqrt{r^2}}{\sqrt{2}}=y\]

OpenStudy (anonymous):

\[\frac{\sqrt{2}\sqrt{r^2}}{2}\]

OpenStudy (anonymous):

where r is your radius

OpenStudy (anonymous):

sorry the last number was 2. i guess the radius is root of 2.